OpenStudy (anonymous):

Mason and Nora decided to swim across the river. Mason began swimming 33 seconds earlier than Nora. Mason swam at a speed of 3 feet per second. Nora swam at a speed of 4 feet per second. For how many seconds had Mason been swimming at the moment when the two swimmers had swam exactly the same distance?

OpenStudy (anonymous):

i forget what the formula is to this

OpenStudy (anonymous):

i have no clue.

OpenStudy (anonymous):

yeah i took something like this a year ago but it totally slipped my mind let me look. what is the lesson you are on?

OpenStudy (anonymous):

so that i know like what to search

OpenStudy (anonymous):

@Starbucks

OpenStudy (anonymous):

its algebra , but like im in the end so i dont really know..

OpenStudy (anonymous):

ok well what is it asking for?

OpenStudy (anonymous):

For how many seconds had Mason been swimming at the moment when the two swimmers had swam exactly the same distance?

OpenStudy (anonymous):

i know that nora caught up to mason at 120 seconds, but that was wrong.

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

that guy i bet can help

OpenStudy (anonymous):

i hoope.

OpenStudy (anonymous):

@deliciotron

OpenStudy (anonymous):

@chosenmatt

OpenStudy (deliciotron):

speed = distance / time but once we have the equation, we can use any of its variations, speed = distance / time distance = speed * time time = distance / speed

OpenStudy (anonymous):

that's it. that's the equation

OpenStudy (deliciotron):

For example, suppose we drive for 2 hours at 30 miles per hour, for a total of 60 miles. If we know the time and the speed, we can find the distance: 2 hours * 30 miles/hour = 60 miles If we know the time and the distance, we can find the speed: 60 miles / 2 hours = 30 miles/hour If we know the speed and the distance, we can find the time: 60 miles / (30 miles/hour) = 2 hours It's a little like having a family of multiplication facts, e.g., 12 = 3 * 4 3 = 12 / 4 4 = 12 / 3

OpenStudy (anonymous):

i dont get how to use that for my question though .

OpenStudy (deliciotron):

Just forgot. lemme think for a sec

OpenStudy (deliciotron):

Does it tell you howw long it took them to be at the same place?

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

i got that myself though, it took 120 seconds.

OpenStudy (deliciotron):

Ok, so 120 seconds times 4 feet per second and 120 seconds times 3 feet per second. That would be the next step.

OpenStudy (deliciotron):

Actually it would be 4 feet per second * 120 second and the same for 3 feeet per second

OpenStudy (deliciotron):

feet*

OpenStudy (deliciotron):

@kobeni-chan

OpenStudy (deliciotron):

Then you see how long it takes for nora to pass mason

OpenStudy (deliciotron):

Did that help?

OpenStudy (anonymous):

no .